Roots Of x 4 +4x 3 -4x-13=0?

Roots Of x4+4x3-4x-13=0?

Grade:

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question

x^4 + 4x^3 - 4x - 13 = 0
x^4 + 4x^3 - 4x - 13 - 3 = -3
x^4 + 4x^3 - 4x - 16 = -3
x^3(x + 4) - 4(x + 4) = -3
(x^3 - 4)(x + 4) = -3
\Rightarrow x = -3.95, 1.512
\Rightarrow x = -0.78\pm 1.25i